media%2Ffdf%2Ffdfa9fe8-7cc4-4e60-8cc8-d4


calculate the amount of oxygen provided to an aerobic culture, expressed as mg of O2/Liz of culture-hour, if 20 L of air, at stp conditions, is bubbled per hour into the culture brother volume of one liter. Assume that only 25% of the oxygen in the bubbled air ends up dissolved in the water, due to mass transfer limitations. Show the calculations you performed to arrive to your answer.
Answer #1

20l of air bubbles /hr into 1 liter broth

Air contain 20% pure oxygen

20liter air X 20/100 = 4 liter oxygen pure oxygen provided

25% is resolved of O2 disolve

Means 4L O2 x25/100=1lit

1liter 02 disolve /liter of culture / hour

1 liter = 1000ml

1000ml of oxygen resolve /liter of culture / hour

ml/L is equivalent to part pr thousand

But the volume of gases is not equivalents to mass it can be calculated by gas law

PV=nRT

where

P

= pressure in atm (1 atm = 1000mb = 101.3 kPa)

V

= volume in liters

n

= moles of gas

R

= gas constant (for these units, =0.0821 L*atm/mol*K

T

= temperature in Kelvins ( T(K) = T(degC) + 273K )

1atmX1lit of oxygen=nX0.0821x293k

n= 1atmx1liter of oxygen/0.0821X293k

356.881molO2/liter

MW of oxygen 32.0g/mole

1 mol contain 32 g oxygen

356.881 mol contain 32x356.881

=11420.19 or 11421 g

1g =1000mg

11421x1000=

Disolve oxygen =11421x10 -3mg O2/liter of culture / hour

Provided oxygen 11421x10-3 x 4=45684 x 10-3mg 02/liter of culture / hour